OpenStudy (anonymous):

Factor expression: 2X^(3/2)+6X^(1/2)-4X^(-1/2)

OpenStudy (anonymous):

\[2x ^{\frac{ 3 }{ 2 }} + 6x ^{\frac{ 1 }{ 2 }} - 4x ^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

2x3=6 2x2=4 2x1=2 ok?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so the factor is: \[2x ^{\frac{ 1 }{ 2 }}(1+3-2^{-1})\]

OpenStudy (anonymous):

wait , in the bracket, isn't it (x+3-....)

OpenStudy (anonymous):

because we factor out x^1/2 from X^3/2

OpenStudy (anonymous):

wait I think I made mistake...

OpenStudy (anonymous):

It will be:

OpenStudy (anonymous):

\[2x ^{\frac{ 1 }{ 2 }}(x ^{3}+3-2^{-1})\]

OpenStudy (anonymous):

ok?got it?

OpenStudy (anonymous):

why is it X^3?

OpenStudy (anonymous):

3/2 - 1/2 = 1

OpenStudy (anonymous):

shouldn't it be just x

OpenStudy (anonymous):

x^1 = x

OpenStudy (anonymous):

you are quite right sorry,sorry,sorry...so please write it.

OpenStudy (anonymous):

but i dont know about the last one ...

OpenStudy (anonymous):

1+1/2 = 3/2

OpenStudy (anonymous):

4x^(-1/2) factor out 2x^1/2, leaves what?

OpenStudy (anonymous):

you are right ;)

OpenStudy (anonymous):

leaves 2^-1?

OpenStudy (anonymous):

wait 'm confused :D let me doing it...

OpenStudy (anonymous):

or 2X^-1

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

ok i think it's like......\[2x ^{1/2} (x+3-\frac{ 2 }{ x })\]

OpenStudy (anonymous):

because 2/x = \[2X ^{-1}\]

OpenStudy (anonymous):

well i'm guessing......

OpenStudy (anonymous):

true :)

OpenStudy (anonymous):

ok cool